OpenStudy (anonymous):

For what values of x does x^2+10x+2=0?

OpenStudy (anonymous):

use the quadratic equation. do you know it?

OpenStudy (anonymous):

yea... I tried it though and still didnt get the right answer! Need help!

OpenStudy (anonymous):

\[-b \pm \sqrt{b^2-(4)(a)(c)}/2a\]

OpenStudy (anonymous):

\[-10\pm \sqrt{100-8}/2\]

OpenStudy (anonymous):

\[-10\pm \sqrt{92}/2\]

OpenStudy (anonymous):

can you finish from here?

OpenStudy (anonymous):

yea! thanks!